Paddled For 3 Hours With A 7 Km/h Current To Reach A Campsite. The Return Trip Against The Current Took 10 Hours. Find The Speed Of The Boat In Still Water?
Answer by mananth(16949)

Paddled For 3 Hours With A 7 Km/h Current To Reach A Campsite. The Return Trip Against The Current Took 10 Hours. Find The Speed Of The Boat In Still Water?
...
speed of boat in still water =x
..
with current speed = x+7: time = 3 hours
against , speed = x-7: time = 10 hours
...
distance = speed * time
3(x+7)=10(x-7)
3x+21=10x-70
7x=91
x=13 km/h speed of boat
...
m.ananth@hotmail.ca